OpenStudy (anonymous):

Simplify negative 9 over 6 divided by 3 over negative 2..

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@UsukiDoll

jimthompson5910 (jim_thompson5910):

hint: flip the second fraction and multiply \[\Large \frac{-9}{6}\div \frac{3}{-2} = \frac{-9}{6}\times \frac{-2}{3}\]

OpenStudy (anonymous):

wht u mean

OpenStudy (anonymous):

???

jimthompson5910 (jim_thompson5910):

did you see how \(\LARGE \frac{3}{-2}\) flipped to \(\LARGE \frac{-2}{3}\) ?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

now multiply straight across and reduce the fraction

OpenStudy (anonymous):

3/2 x 2/3

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

this here \[\Large \frac{-9}{6}\times \frac{-2}{3}\]

OpenStudy (anonymous):

1

jimthompson5910 (jim_thompson5910):

what did you get before you got 1

OpenStudy (anonymous):

is 1 right???

jimthompson5910 (jim_thompson5910):

I'm just checking what you got before you got 1. Before you reduced the fraction.

OpenStudy (anonymous):

i guessed sry

jimthompson5910 (jim_thompson5910):